1. Evaluating the results for accuracy

In an experiment, the concentration of a protein in a solution was measured by a student to be 273µg/ml, but the true value of the concentration was 260µg/ml. What is the percent error in this measurement?

Relative error: (2 marks)

Comment on the accuracy: (2 marks)

2. Examine the following data set for the presence of outliers using the Q-test.(4 marks)

Replicate # Values

1 0.200


2 0.186


3 0.110


4 0.196


5 0.193


1
Expert's answer
2021-11-19T08:28:01-0500

273-260= 13

= 13/260×100= 5% error